# Date Fragments: A Hidden Bottleneck of Tokenisation for Temporal Reasoning

### v2

@@ Abstract
Modern BPE tokenisers often split calendar dates into meaningless fragments, e.g., “20250312” $\rightarrow$ “202”, “503”, “12”, inflating token counts and obscuring the inherent structure needed for robust temporal reasoning. In this work, we (1) introduce a simple yet interpretable metric, termed date fragmentation ratio, that measures how faithfully a tokeniser preserves multi-digit date components; (2) release DateAugBench, a suite of 6500 examples spanning three temporal reasoning tasks: context-based date resolution, format-invariance puzzles, and date arithmetic across historical, contemporary, and future time periods; and (3) through layer-wise probing and causal attention-hop analyses, uncover an emergent date-abstraction mechanism whereby large language models stitch together the fragments of month, day, and year components for temporal reasoning. Our experiments show that excessive fragmentation correlates with accuracy drops of up to 10 points on uncommon dates like historical and futuristic dates. Further, we find that the larger the model, the faster the emergent date abstraction heals date fragments. Lastly, we observe a reasoning path that LLMs follow to assemble date fragments, typically differing from human interpretation (year $\rightarrow$ month $\rightarrow$ day). Our datasets and code are made publicly available here
@@ 1 Introduction
Understanding and manipulating dates is a deceptively complex challenge for modern large language models (LLMs). Unlike ordinary words, dates combine numeric and lexical elements in rigidly defined patterns—ranging from compact eight-digit strings such as 20250314 to more verbose forms like “March 14, 2025” or locale-specific variants such as “14/03/2025.” Yet despite their structured nature, these date expressions often fall prey to subword tokenisers that fragment them into semantically meaningless pieces. A tokeniser that splits “2025-03-14” into “20”, “25”, “-0”, “3”, “-1”, “4” not only inflates the token count but also severs the natural boundaries of year, month, and day. This fragmentation obscures temporal cues and introduces a hidden bottleneck: even state-of-the-art LLMs struggle to resolve, compare, or compute dates accurately when their internal representations have been so badly fragmented. This issue has a critical impact on real-world applications:
@@
Mis-tokenised dates can undermine scheduling and planning workflows, leading to erroneous calendar invites or appointments Vasileiou and Yeoh (2024). They can skew forecasting models in domains ranging from time-series analysis Tan et al. (2024); Chang et al. (2023) to temporal knowledge graph reasoning Wang et al. (2024). In digital humanities and historical scholarship, incorrect splitting of date expressions may corrupt timelines and misguide interpretative analyses Zeng (2024). As LLMs are increasingly deployed in cross-temporal applications, such as climate projection Wang and Karimi (2024), economic forecasting Carriero et al. (2024); Bhatia et al. (2024), and automated curriculum scheduling Vasileiou and Yeoh (2024), the brittleness introduced by subword fragmentation poses a risk of propagating temporal biases and inaccuracies into downstream scientific discoveries and decision-making systems Tan et al. (2024).
@@ 3 DateAugBench
Dataset and Task # Formats # Raw Size Evaluation
Example GT
Context based 6 500 3000 Which team did Omid Namazi play for in 06/10/1990? Maryland Bays
Date Format Switching 10 150 1500 Are 20251403 and March 14th 2025 referring to the same date? Yes
Date Arithmetic 5 400 2000 What date is 10,000 days before 5/4/2025? 18 November 1997; 17 December 1997
Total 21 1500 6500
@@ 4.1 Date Tokenisation
$F$ $=0.10\times\mathbf{1}_{\mathrm{split}}+0.10\times\mathbf{1}_{\mathrm{delimiter}}$ (2)
$\quad+0.05\times\bigl{(}N-N_{b}\bigr{)}+0.30\times\theta$
This date fragmentation ratio is pivotal because tokenisation inconsistencies directly impair a model’s ability to represent and reason over temporal inputs. When date strings are split non-intuitively, models face inflated token sequences and fragmented semantic cues, potentially leading to errors in tasks such as chronological comparison, date arithmetic, and context-based resolution.
@@ 4.3 Internal Representations
We use four Qwen2.5 Yang et al. (2024) model checkpoints (0.5B, 1.5B, 3B, and 7B parameters) to trace how temporal information is processed internally across different layers. During inference, each question is prefixed with a fixed system prompt and a chain-of-thought cue, then passed through the model in evaluation mode. At each layer $i$, we extract the hidden-state vector corresponding to the final token position, yielding an embedding $h_{i}\in\mathbb{R}^{d}$ for that layer. Repeating over all examples produces a collection of layer-wise representations for positive and negative cases. We then quantify the emergence of temporal reasoning by training lightweight linear probes on these embeddings. For layer $i$, the probe is trained to distinguish “same-date” (positive) vs “different-date” (negative) examples. To explain when the model’s date understanding is achieved, we define the tokenisation compensation point as the layer at which the model’s representation correctly represents the date in the given prompt. We experiment with this idea across various model sizes, aiming to test our hypothesis: larger models would recover calendar-level semantics from fragmented tokens at earlier stages, i.e., tokenisation compensation is accomplished at early layers, as illustrated in Figure 2.
@@
We introduce a framework intended to understand in which order date fragments are stitched together for LLMs to answer a temporal question. Figure 1 depicts the idea of our framework: given an input prompt requiring a date resolution (e.g., “Is 28052025 the same date as 28th of May 2025?”), we define two sets of tokens: (1) concept tokens corresponding to year, month, and day fragments, and (2) decision tokens corresponding to the model answer (“yes” or “no”). Our framework aims to identify a stitching path for temporal reasoning, or reasoning path for short. A reasoning path is defined as a sequence of tokens containing date fragments and the model answer [^1] The idea of reasoning paths was introduced by Lindsey et al. (2025), which we leverage to interpret how LLMs address date fragments for temporal reasoning. . Given that there are multiple potential paths, we score each path and select the highest-scoring one as the LLM’s reasoning path for the given prompt. To score a reasoning path, our idea is the following: we identify when a date fragment or model answer is activated, by which input token and at which layer, and then determine how important each input token is for the date fragment and model answer. Our idea is implemented by using two different approaches: (i) next token prediction (§A.2.1): how likely a date fragment and model answer follows a given input token and (ii) token importance (§A.2.2): how important an input token is to a date fragment and model answer (by replacing the input token with a random token). Lastly, we combine the results of the two approaches to yield the final score of a reasoning path (§A.2.3). This causal framework not only pinpoints where and when date fragments are activated, but also in which order they are stitched together to yield the model answer.
@@ 5.1 Date fragmentation
Table 2 reports the mean date fragmentation ratio across four time periodsPast (pre–2000), Near Past (2000–2009), Present (2010–2025), and Future (post–2025)—for each evaluated model. A ratio of 0.00 signifies perfect alignment with our rule-based baseline tokeniser, whereas higher values indicate progressively greater fragmentation. The rule-based Baseline unsurprisingly attains the maximal ratio of 0.00 in all periods, serving as a lower bound. Among neural architectures, OLMo Groeneveld et al. (2024) demonstrates the highest robustness, with an average fragmentation ratio of 0.15, closely followed by GPT-3 at 0.16. Both maintain strong fidelity across temporal splits, although performance dips modestly in the Future category (0.25), reflecting novel token sequences not seen during pre-training.
@@
A closer look, from Table 4, at sub-token granularity further explains these trends. Llama 3 Touvron et al. (2023) and the GPT OpenAI et al. (2023) families typically segment each date component into three-digit sub-tokens (e.g., “202”, “504”, “03”), thus preserving the semantic unit of “MMDDYYYY” as compact pieces. OLMo Groeneveld et al. (2024) splits the date tokens into two digit tokens (e.g., “20”, “25”). By contrast, Qwen Yang et al. (2024) and Gemma Team et al. (2024) models break dates into single-digit tokens (e.g., “2”, “5”), whereas Phi Abdin et al. (2024) divides it into single-digit tokens with an initial token (e.g. “_”, “2”, “0”, “2”, “5”), inflating the token count. Although single-digit tokenisation can enhance models’ ability to perform arbitrary numeric manipulations (by treating each digit as an independent unit), it comes at the expense of temporal abstraction: the tight coupling between day, month, and year is lost, inflating the compression penalty and increasing the $\theta$ divergence from the baseline.
@@ 5.2 DateFragBench Evaluation
Figure 3 plots date fragmentation ratio against resolution accuracy, with 24 data points across six models and four temporal splits. Accuracy rises as we move from Past (1600-2000) to Near Past (2000–2009) and peaks in the Present (2010–2025), mirroring the negative correlation between fragmentation and accuracy (dashed line, Pearson correlation of $-0.61$). We note that the correlation is not particularly strong. This is because (i) for some models (e.g., Phi), their date fragmentation ratios remain unchanged across temporal data splits and (ii) models differ greatly by their sizes: a larger model could outperform a substantially smaller model in terms of temporal reasoning performance, even if the former has a much higher fragmentation ratio.
@@
Figure 4 plots model accuracy against date fragmentation ratio across six date formats and six LLMs. A moderate negative trend emerges (dashed line, Pearson correlation of $-0.42$): formats that contain explicit separators (DD-MM-YYYY, DD/MM/YYYY, YYYY/MM/DD) are tokenised into more pieces and, in turn, resolved more accurately than compact, separator-free strings (DDMMYYYY, MMDDYYYY, YYYYMMDD). As shown in Table 6, GPT-4o-mini tops every format and receives a moderate performance drop from 71.2 % on DD/MM/YYYY to 61.2 % on DDMMYYYY, with the highest overall average (66.3 %). OLMo-2-7B and Qwen-2.5-14B both exceed 70 % on the highly fragmented YYYY/MM/DD form, but slip into the low 50s on MMDDYYYY and YYYYMMDD. Lower date fragmentation ratio models, such as Llama-3.1-8B and Phi-3.5, lag behind; their accuracy plunges below 40 %. Even so, all models score much better on separator-rich formats compared to the date formats without separators. In summary, model accuracy is correlated to how cleanly a model can tokenise the string into interpretable tokens: more visual structure (slashes or dashes) means lower fragmentation, which suggests more straightforward reasoning, and in turn, leads to better performance.
@@ 6 In which layer do LLMs compensate for date fragmentation?
To pinpoint in which layer a model learns to recognize two equivalent dates, we define the tokenisation compensation point (TCP) as the earliest layer at which a lightweight linear probe on the hidden state achieves above-chance accuracy, which is defined as 80%, on the date equivalence task. Figure 5(a) reports TCPs for the Dates_Past benchmark (1600–2010): Qwen2.5-0.5B reaches TCP at layer 12 (50% depth), Qwen2.5-1.5B at layer 15 (53.6%), Qwen2.5-3B at layer 8 (22.2%), and Qwen2.5-7B at layer 4 (14.3%). The leftward shift of the 3B and 7B curves suggests how larger models recover calendar-level semantics from fragmented tokens more rapidly. Figure 5(b) shows the Dates_Present benchmark (2010–2025), where only the 1.5B, 3B, and 7B models surpass TCP—at layers 16 (57.1%), 21 (58.3%), and 17 (60.7%), respectively—while the 0.5B model never does. The deeper TCPs here reflect extra layers needed to recombine the two-digit “20” prefix, which is fragmented unevenly by the tokeniser. In Figure 10, we evaluate Dates_Future (2025–2599), where novel four-digit sequences exacerbate fragmentation. Remarkably, TCPs mirror the Past regime: layers 12, 15, 8, and 4 for the 0.5B, 1.5B, 3B, and 7B models, respectively. This parallelism indicates that model scale dictates how quickly LLMs can compensate for date fragmentation to achieve high accuracy, even when dates are novel.
@@ 7 How do LLMs stitch date fragments for temporal reasoning?
To investigate how LLMs like Llama 3 Touvron et al. (2023) internally stitch date fragments to yield a model answer, we apply our casual framework to identify the model’s reasoning path over a specific prompt. Figure 6 plots model layers on the $y$ axis against prompt tokens (e.g., Is 03122025 a valid date?) on the $x$ axis. Green arrows mark the reasoning path with the highest score that is responsible for generating the answer “yes”. Date fragments “25”, “220”, “031”, and the model answer “yes” are activated in sequence at layer 26-27 by the input tokens “is”, “031”, “a” and “Answer” respectively. As such, the model performs a kind of discrete, step-by-step token aggregation, stitching together substrings of the input until a binary valid/invalid verdict emerges.
@@
In contrast, human readers parse dates by immediately mapping each component to a coherent temporal schema: “03” is March, “12” is day of month, “2025” is year, and then checking whether the day falls within the calendar bounds of that month. Humans bring rich world knowledge of calendars and leap-year rules to bear in parallel. However, LLMs exhibit no explicit calendar “module”; instead, they rely on learned statistical associations between digit-patterns and the training-time supervisory signal for “valid date”. The reasoning path in Figure 6 thus illustrates a fundamentally different mechanism of date comprehension in LLMs, based on date fragments re-routing rather than holistic semantic interpretation. We repeated causal tracing on 100 date strings in 6 different date formats to test whether the reasoning path difference between human and LLMs is consistent across date formats. In most of cases, we observe that model reasoning paths are not aligned with human interpretation (year $\rightarrow$ month $\rightarrow$ day), rather rely on sub-word fragments that statistically represent year, month, and day, and stitch these date fragments in a flexible order that is subject to date formats (see examples in Figures 7-8). However, such a reasoning path becomes tricky when a date is greatly fragmented: given the date abstraction is learned from frequency rather than hard-coded rules, the abstraction is biased toward standard Western formats and contemporary years. As a result, a model often addresses popular dates (in the same format) with similar reasoning paths. However, the reasoning path becomes obscure on rare, historical, or locale-specific strings outside the distribution of pre-training data (see Figure 9).

### v3

@@ Abstract
Modern BPE tokenisers often split calendar dates into meaningless fragments, e.g., “20250312” $\rightarrow$ “202”, “503”, “12”, inflating token counts and obscuring the inherent structure needed for robust temporal reasoning. In this work, we (1) introduce a simple yet interpretable metric, termed date fragmentation ratio, that measures how faithfully a tokeniser preserves multi-digit date components; (2) release DateAugBench, a suite of 6500 examples spanning three temporal reasoning tasks: context-based date resolution, format-invariance puzzles, and date arithmetic across historical, contemporary, and future time periods; and (3) through layer-wise probing and causal attention-hop analyses, uncover an emergent date-abstraction mechanism whereby large language models stitch together the fragments of month, day, and year components for temporal reasoning. Our experiments show that excessive fragmentation correlates with accuracy drops of up to 10 points on uncommon dates like historical and futuristic dates. Further, we find that the larger the model, the faster the emergent date abstraction heals date fragments. Lastly, we observe a reasoning path that LLMs follow to assemble date fragments, typically differing from human interpretation (year $\rightarrow$ month $\rightarrow$ day). Our datasets and code are made publicly available here.
@@ 1 Introduction
Understanding and manipulating dates is a deceptively complex challenge for modern large language models (LLMs). Unlike ordinary words, dates combine numeric and lexical elements in rigidly defined patterns—ranging from compact eight‐digit strings such as 20250314 to more verbose forms like “March 14, 2025” or locale‐specific variants such as “14/03/2025.” Yet despite their structured nature, these date expressions often fall prey to subword tokenisers that fragment them into semantically meaningless pieces. A tokeniser that splits “2025-03-14” into “20”, “25”, “-0”, “3”, “-1”, “4” not only inflates the token count but also severs the natural boundaries of year, month, and day. This fragmentation obscures temporal cues and introduces a hidden bottleneck: even state-of-the-art LLMs struggle to resolve, compare, or compute dates accurately when their internal representations have been so badly fragmented. This issue has a critical impact on real-world applications:
@@
Mis‐tokenised dates can undermine scheduling and planning workflows, leading to erroneous calendar invites or appointments Vasileiou and Yeoh (2024). They can skew forecasting models in domains ranging from time‐series analysis Tan et al. (2024); Chang et al. (2023) to temporal knowledge graph reasoning Wang et al. (2024). In digital humanities and historical scholarship, incorrect splitting of date expressions may corrupt timelines and misguide interpretative analyses Zeng (2024). As LLMs are increasingly deployed in cross‐temporal applications, such as climate projection Wang and Karimi (2024), economic forecasting Carriero et al. (2024); Bhatia et al. (2024), and automated curriculum scheduling Vasileiou and Yeoh (2024), the brittleness introduced by subword fragmentation poses a risk of propagating temporal biases and inaccuracies into downstream scientific discoveries and decision‐making systems Tan et al. (2024).
@@ 3 DateAugBench
2 Dataset and Task # Formats # Raw Size Evaluation Example GT Context based 6 500 3000 Which team did Omid Namazi play for in 06/10/1990? Maryland Bays Date Format Switching 10 150 1500 Are 20251403 and March 14th 2025 referring to the same date? Yes Date Arithmetic 5 400 2000 What date is 10,000 days before 5/4/2025? 18 November 1997; 17 December 1997 Total 21 1500 6500
@@ 4.1 Date Tokenisation
[…]
This date fragmentation ratio is pivotal because tokenisation inconsistencies directly impair a model’s ability to represent and reason over temporal inputs. When date strings are split non-intuitively, models encounter inflated token sequences and fragmented semantic cues, which can potentially lead to errors in tasks such as chronological comparison, date arithmetic, and context-based resolution.
Validation of Date Fragmentation Ratio.
To ensure our custom metric is well-founded, we performed a two-part validation. First, we conducted a human evaluation study, in which we found that our F metric’s scores align strongly with human judgments of "fragmentation severity" (Spearman’s $\rho=0.84$), significantly outperforming standard metrics like BLEU ($\rho=0.52$). Second, we used a data-driven approach to learn the metric’s weights by training a model to predict the human severity scores from our fragmentation components. This process confirmed that our intuitively chosen weights accurately reflect the factors driving human perception of fragmentation. For a detailed breakdown of the human evaluation protocol and the data-driven validation, please see Appendix A.3.
@@ 4.3 Internal Representations
We use four Qwen2.5 Yang et al. (2024) model checkpoints (0.5B, 1.5B, 3B, and 7B parameters) to trace how temporal information is processed internally across different layers. During inference, each question is prefixed with a fixed system prompt and a chain‐of‐thought cue, then passed through the model in evaluation mode. At each layer $i$, we extract the hidden‐state vector corresponding to the final token position, yielding an embedding $h_{i}\in\mathbb{R}^{d}$ for that layer. Repeating over all examples produces a collection of layer‐wise representations for positive and negative cases. We then quantify the emergence of temporal reasoning by training lightweight linear probes on these embeddings. For layer $i$, the probe is trained to distinguish “same‐date” (positive) vs “different‐date” (negative) examples. To explain when the model’s date understanding is achieved, we define the tokenisation compensation point as the layer at which the model’s representation correctly represents the date in the given prompt. We experiment with this idea across various model sizes, aiming to test our hypothesis: larger models would recover calendar‐level semantics from fragmented tokens at earlier stages, i.e., tokenisation compensation is accomplished at early layers, as illustrated in Figure 2.
@@
We introduce a framework intended to understand in which order date fragments are stitched together for LLMs to answer a temporal question. Figure 1 depicts the idea of our framework: given an input prompt requiring a date resolution (e.g., “Is 28052025 the same date as 28th of May 2025?”), we define two sets of tokens: (1) concept tokens corresponding to year, month, and day fragments, and (2) decision tokens corresponding to the model answer (“yes” or “no”). Our framework aims to identify a stitching path for temporal reasoning, or reasoning path for short. A reasoning path is defined as a sequence of tokens containing date fragments and the model answer [^1] The idea of reasoning paths was introduced by Lindsey et al. (2025), which we leverage to interpret how LLMs address date fragments for temporal reasoning. . Given that there are multiple potential paths, we score each path and select the highest‐scoring one as the LLM’s reasoning path for the given prompt. To score a reasoning path, our idea is the following: we identify when a date fragment or model answer is activated, by which input token and at which layer, and then determine how important each input token is for the date fragment and model answer. Our idea is implemented by using two different approaches: (i) next token prediction (§A.2.1): how likely a date fragment and model answer follows a given input token and (ii) token importance (§A.2.2): how important an input token is to a date fragment and model answer (by replacing the input token with a random token). Lastly, we combine the results of the two approaches to yield the final score of a reasoning path (§A.2.3). This causal framework not only pinpoints where and when date fragments are activated, but also in which order they are stitched together to yield the model answer.
@@ 5.1 Date fragmentation
Table 2 reports the mean date fragmentation ratio across four time periodsPast (pre–2000), Near Past (2000–2009), Present (2010–2025), and Future (post–2025)—for each evaluated model. A ratio of 0.00 signifies perfect alignment with our rule‐based baseline tokeniser, whereas higher values indicate progressively greater fragmentation. The rule‐based Baseline unsurprisingly attains the maximal ratio of 0.00 in all periods, serving as a lower bound. Among neural architectures, OLMo Groeneveld et al. (2024) demonstrates the highest robustness, with an average fragmentation ratio of 0.15, closely followed by GPT-3 at 0.16. Both maintain strong fidelity across temporal splits, although performance dips modestly in the Future category (0.25), reflecting novel token sequences not seen during pre-training.
@@
A closer look, from Table 4, at sub-token granularity further explains these trends. Llama 3 Touvron et al. (2023) and the GPT OpenAI et al. (2023) families typically segment each date component into three-digit sub-tokens (e.g., “202”, “504”, “03”), thus preserving the semantic unit of “MMDDYYYY” as compact pieces. OLMo Groeneveld et al. (2024) splits the date tokens into two digit tokens (e.g., “20”, “25”). By contrast, Qwen Yang et al. (2024) and Gemma Team et al. (2024) models break dates into single-digit tokens (e.g., “2”, “5”), whereas Phi Abdin et al. (2024) divides it into single-digit tokens with an initial token (e.g. “_”, “2”, “0”, “2”, “5”), inflating the token count. Although single‐digit tokenisation can enhance models’ ability to perform arbitrary numeric manipulations (by treating each digit as an independent unit), it comes at the expense of temporal abstraction: the tight coupling between day, month, and year is lost, inflating the compression penalty and increasing the $\theta$ divergence from the baseline.
@@ 5.2 DateFragBench Evaluation
Figure 3 plots date fragmentation ratio against resolution accuracy, with 24 data points across six models and four temporal splits. Accuracy rises as we move from Past (1600-2000) to Near Past (2000–2009) and peaks in the Present (2010–2025), mirroring the negative correlation between fragmentation and accuracy (dashed line, Pearson correlation of 0.61-0.61). We note that the correlation is not particularly strong. This is because (i) for some models (e.g., Phi), their date fragmentation ratios remain unchanged across temporal data splits and (ii) models differ greatly by their sizes: a larger model could outperform a substantially smaller model in terms of temporal reasoning performance, even if the former has a much higher fragmentation ratio.
@@
Figure 4 plots model accuracy against date fragmentation ratio across six date formats and six LLMs. A moderate negative trend emerges (dashed line, Pearson correlation of 0.42-0.42): formats that contain explicit separators (DD-MM-YYYY, DD/MM/YYYY, YYYY/MM/DD) are tokenised into more pieces and, in turn, resolved more accurately than compact, separator-free strings (DDMMYYYY, MMDDYYYY, YYYYMMDD). As shown in Table 9, GPT-4o-mini tops every format and receives a moderate performance drop from 71.2 % on DD/MM/YYYY to 61.2 % on DDMMYYYY, with the highest overall average (66.3 %). OLMo-2-7B and Qwen-2.5-14B both exceed 70 % on the highly fragmented YYYY/MM/DD form, but slip into the low 50s on MMDDYYYY and YYYYMMDD. Lower date fragmentation ratio models, such as Llama-3.1-8B and Phi-3.5, lag behind; their accuracy plunges below 40 %. Even so, all models score much better on separator-rich formats compared to the date formats without separators. In summary, model accuracy is correlated to how cleanly a model can tokenise the string into interpretable tokens: more visual structure (slashes or dashes) means lower fragmentation, which suggests more straightforward reasoning, and in turn, leads to better performance.
@@ 6 In which layer do LLMs compensate for date fragmentation?
To pinpoint in which layer a model learns to recognize two equivalent dates, we define the tokenisation compensation point (TCP) as the earliest layer at which a lightweight linear probe on the hidden state achieves above‐chance accuracy, which is defined as 80%, on the date equivalence task. Figure 5(a) reports TCPs for the Dates_Past benchmark (1600–2010): Qwen2.5-0.5B reaches TCP at layer 12 (50% depth), Qwen2.5-1.5B at layer 15 (53.6%), Qwen2.5-3B at layer 8 (22.2%), and Qwen2.5-7B at layer 4 (14.3%). The leftward shift of the 3B and 7B curves suggests how larger models recover calendar‐level semantics from fragmented tokens more rapidly. Figure 5(b) shows the Dates_Present benchmark (2010–2025), where only the 1.5B, 3B, and 7B models surpass TCP—at layers 16 (57.1%), 21 (58.3%), and 17 (60.7%), respectively—while the 0.5B model never does. The deeper TCPs here reflect extra layers needed to recombine the two‐digit “20” prefix, which is fragmented unevenly by the tokeniser. In Figure 10, we evaluate Dates_Future (2025–2599), where novel four‐digit sequences exacerbate fragmentation. Remarkably, TCPs mirror the Past regime: layers 12, 15, 8, and 4 for the 0.5B, 1.5B, 3B, and 7B models, respectively. This parallelism indicates that model scale dictates how quickly LLMs can compensate for date fragmentation to achieve high accuracy, even when dates are novel.
@@ 7 How do LLMs stitch date fragments for temporal reasoning?
To investigate how LLMs like Llama 3 Touvron et al. (2023) internally stitch date fragments to yield a model answer, we apply our casual framework to identify the model’s reasoning path over a specific prompt. Figure 6 plots model layers on the $y$ axis against prompt tokens (e.g., Is 03122025 a valid date?) on the $x$ axis. Green arrows mark the reasoning path with the highest score that is responsible for generating the answer “yes”. Date fragments “25”, “220”, “031”, and the model answer “yes” are activated in sequence at layer 26-27 by the input tokens “is”, “031”, “a” and “Answer” respectively. As such, the model performs a kind of discrete, step‐by‐step token aggregation, stitching together substrings of the input until a binary valid/invalid verdict emerges.
@@
In contrast, human readers parse dates by immediately mapping each component to a coherent temporal schema: “03” is March, “12” is day of month, “2025” is year, and then checking whether the day falls within the calendar bounds of that month. Humans bring rich world knowledge of calendars and leap‐year rules to bear in parallel. However, LLMs exhibit no explicit calendar “module”; instead, they rely on learned statistical associations between digit‐patterns and the training‐time supervisory signal for “valid date”. The reasoning path in Figure 6 thus illustrates a fundamentally different mechanism of date comprehension in LLMs, based on date fragments re‐routing rather than holistic semantic interpretation. We repeated causal tracing on 100 date strings in 6 different date formats to test whether the reasoning path difference between human and LLMs is consistent across date formats. In most of cases, we observe that model reasoning paths are not aligned with human interpretation (year $\rightarrow$ month $\rightarrow$ day), rather rely on sub-word fragments that statistically represent year, month, and day, and stitch these date fragments in a flexible order that is subject to date formats (see examples in Figures 7-8). However, such a reasoning path becomes tricky when a date is greatly fragmented: given the date abstraction is learned from frequency rather than hard-coded rules, the abstraction is biased toward standard Western formats and contemporary years. As a result, a model often addresses popular dates (in the same format) with similar reasoning paths. However, the reasoning path becomes obscure on rare, historical, or locale-specific strings outside the distribution of pre-training data (see Figure 9).
8 Discussion
The moderate Pearson correlations of -0.61 (by temporal regime) and -0.42 (by date format) are a significant finding in themselves. They confirm that date fragmentation is a consistent and independent bottleneck, while also highlighting that it is not the sole factor influencing performance. The remaining variance is naturally explained by confounding factors such as model architecture, scale, and pre-training data exposure. For instance, a larger model may have a greater capacity to "heal" a poorly tokenised date, partially masking the negative impact of a high fragmentation ratio. Nonetheless, our results demonstrate that, all else being equal, higher fragmentation consistently predicts a drop in accuracy. This reveals a fundamental impediment that exists at the input level, before the model’s core reasoning layers are even engaged.
Our findings also shed light on the debate between memorisation and actual logical reasoning in LLMs. The performance disparity between "Present" dates and "Past" or "Future" dates (Figure 3) suggests that models heavily rely on statistical patterns and memorised facts from their pre-training data. For common contemporary dates, strong learned associations allow models to effectively parse and reason about them, even if tokenisation is suboptimal. However, for less frequent historical or novel future dates, this reliance on memorisation becomes a liability. The "date abstraction" mechanism struggles, and models must generalise from sparser data, leading to the observed accuracy drops. This contrasts with human date parsing, which leverages explicit, rule-based calendar knowledge rather than frequency-based recall.